OpenStudy (anonymous):

Solve for k. 8k + 2m = 3m + k k = 70m k = 7m k = 7/m k = m/7

OpenStudy (anonymous):

which one

OpenStudy (anonymous):

m/7

OpenStudy (anonymous):

So 8k +2m = 3m +k Move all K's on the same side

OpenStudy (anonymous):

So subtract k to make the other side 7k+2m=3m And then move the 2m to the other side by subtracting it.

OpenStudy (anonymous):

7k=m

OpenStudy (anonymous):

So isolate k by dividing by 7

OpenStudy (anonymous):

k= m/7

OpenStudy (anonymous):

so heavenfun13 is right

OpenStudy (anonymous):

Yup. But you should learn how to do it, not just get the answer. It gets much harder from here unfortunately.

OpenStudy (anonymous):

ok